# Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems

### v1

@@ Abstract
Ensuring LLM reliability requires not only solving complex problems but also recognizing when a problem is unsolvable. Current models often struggle to distinguish objective unsolvability (inherent contradictions in the problem) from subjective capability limitations (problems beyond the model’s competence), which leads to hallucinations and overconfidence. To address this, we propose UnsolvableQA and UnsolvableRL to solve feasible problems, detect inherent contradictions, and prudently refuse tasks beyond capability. Specifically, we construct UnsolvableQA, a dataset of paired solvable and unsolvable instances derived via a dual-track methodology: programmatic generation for logic puzzles and a novel “Reverse Construction” method that injects contradictions into valid reasoning chains for mathematics. Building on this dataset, we introduce UnsolvableRL, a reinforcement learning framework with three reward components jointly accounting for accuracy, unsolvability, and difficulty. Empirical results show that our approach achieves near-perfect unsolvability detection while also improving accuracy on solvable tasks. Crucially, we identify Capability Collapse, demonstrating that explicit exposure to unsolvable data is indispensable for preventing models from becoming systematically overconfident. Our code and data are available at https://github.com/sfasfaffa/unsolvableQA.
@@ 1 Introduction
As Large Language Models (LLMs) achieve strong performance on complex reasoning tasks (chen2025towards; wei2022chain; guo2025deepseek; jaech2024openai), a key research direction (qin2025largelanguagemodelsmeet) is ensuring they recognize their own limitations and behave reliably. Earlier work on confidence calibration (xiong2023can; li2024confidence; yoon2025reasoning) and selective classification (fisch2022calibrated) during reasoning underscores the need for models to know when not to answer. Further, studies on model capability boundaries (chen2024unlocking) examine model self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; chen2025aware). More recently, TruthRL (wei2025truthrl) encourages models to refuse when uncertain through a ternary reward mechanism.
@@
Figure 1: Comparison of alignment paradigms. We treat the solution space (solvable vs. unsolvable) and question difficulty (easy vs. hard) as orthogonal dimensions, requiring models to objectively detect contradictions and subjectively calibrate confidence.
@@
However, a key limitation of these approaches is that they mainly focus on refusing questions that are solvable but currently beyond the model’s capabilities (“Need to Calibrate” region in Figure 1 (a)). This focus neglects a complementary and arguably more fundamental challenge: identifying problems that are inherently unsolvable due to logical contradictions. As illustrated by the gray “Be ignored” region in Figure 1 (a), existing models treat such an unsolvability as mere high difficulty or hallucinate answers even for simple contradictions.
@@
To address this, our work distinguishes these scenarios by treating Solution Space (solvable vs. unsolvable) and Question Difficulty (easy vs. hard) as orthogonal dimensions (Figure 1 (b)). We argue that a reliable AI system must exhibit two behaviors: (1) Objective Detection, accurately identifying and flagging internally contradictory problems as “unsolvable” (top-left yellow quadrant); and (2) Subjective Calibration, safely refusing theoretically solvable problems that exceed the model’s current reasoning capacity. Without this distinction, models may confidently pursue hallucinations or unnecessarily reject solvable tasks.
@@
First, to enable deep analysis of unsolvability, we present a fully automated pipeline and introduce UnsolvableQA, a dataset of paired solvable and unsolvable instances spanning logic, puzzles, and mathematics. For programmatically verifiable tasks, we create provably unsolvable instances by systematically injecting logical contradictions at the code level. For domains such as mathematics that lack direct programmatic verification, inspired by lightman2024lets; ling2023deductive, we propose a “Reverse Construction” methodology. This LLM-driven approach starts from a solvable problem and its correct Chain-of-Thought (CoT), inserts a controlled logical contradiction into the CoT, and reverse-engineers a corresponding unsolvable problem aligned with the flawed reasoning.
To train models that remain robust on both solvable and unsolvable inputs, we design UnsolvableRL, a dynamic reward alignment scheme based on GRPO (shao2024deepseekmath) that enables the model to detect inherent unsolvability and calibrate its confidence. Specifically, we assign distinct optimization objectives to two refusal types: (1) inherent unsolvability is formulated as a supervised classification task, whereas (2) capability-based calibration is governed by a dynamic confidence threshold with the model’s evolving performance.
Empirical results show a synergistic effect: UnsolvableRL achieves near-perfect precision in unsolvability detection and substantially higher reasoning accuracy on solvable tasks, nearly doubling the combined score. Our analysis also reveals a capability collapse phenomenon: without explicit exposure to unsolvable data, especially smaller models face a severe decline in self-monitoring, becoming overconfident in their outputs.
@@
  • We propose a comprehensive alignment framework that treats objective unsolvability and subjective capability as orthogonal dimensions, establishing a robust three-way decision boundary for reliable reasoning.
  • We introduce UnsolvableQA, a dataset constructed via a novel “Reverse Construction” methodology. This approach bridges the data gap by systematically injecting logical contradictions into valid Chain-of-Thought traces to generate verifiable unsolvable instances.
  • We propose UnsolvableRL, a training paradigm leveraging dynamic confidence thresholds. This mechanism enables models to simultaneously learn to solve feasible tasks, detect inherent contradictions, and calibrate confidence on difficult problems.
  • Empirical results demonstrate synergistic improvements in both detection and reasoning. Crucially, we identify the phenomenon of Capability Collapse, revealing that explicit unsolvable data is essential to prevent model overconfidence.
@@ 2 Related Work
There are two primary areas related to our work: automated dataset construction for reasoning and reinforcement learning approaches for incentivizing selective refusal.
As the reliance of Large Language Models on data quality increases, Automated Dataset Construction (ADC) has emerged as a prominent research direction (liu2024automatic; wang2023data). In the realm of reasoning, recent efforts primarily focus on ensuring data solvability and verifiability through formal or programmatic means. Representative works such as Enigmata (chen2025enigmata) and BIGGSM (chen2024unlocking) have significantly enriched the resources for solvable reasoning tasks by employing synthetic puzzles and difficulty-graded annotations, respectively. Concurrently, research on verifying the correctness of Chain-of-Thought (CoT) traces has also been advancing (ling2023deductive; chen2025towards). In contrast, progress in dataset construction regarding problem unsolvability has been relatively slow. Although earlier benchmarks like SQuAD 2.0 (rajpurkar2018knowdontknowunanswerable) attempted to include unanswerable questions, their task formats are overly simplistic (primarily extractive span selection) and fail to effectively evaluate the refusal capabilities of modern models in complex logical reasoning scenarios. Consequently, existing data synthesis efforts remain largely concentrated on solvable instances, leaving a significant research gap.
@@ 3 UnsolvableQA Construction
Existing benchmarks mainly target solvable tasks, leaving models prone to hallucinations when judging feasibility. To address this issue, we introduce UnsolvableQA, a dataset for detecting problem inherent contradictions.
To achieve this, we use a dual-track methodology (illustrated in Figure 2) tailored to the specific verification constraints of different domains.
@@ 3.1 Mathematics Problem Generation via Reverse Construction
Data Generation Process: First, we validate each seed problem by confirming that a strong reasoning model (DeepSeek-Reasoner) can correctly solve it and produce a complete chain-of-thought, ensuring sufficient reasoning capability. Then, we employ DeepSeek-Reasoner to design logical contradictions via two complementary strategies: Constraint Contradiction, which introduces conflicting external conditions, and Axiom Contradiction, which modifies reasoning steps to violate fundamental mathematical axioms. The model generates a structured plan specifying the location and mechanism of contradiction injection.
Validation Process: Based on this plan, the model then reverse-engineers a new problem statement that preserves the stylistic and complexity characteristics of the seed problem while embedding the designed contradiction. Finally, we implement a two-tier verification protocol: the generated problem is first presented to DeepSeek-Reasoner to autonomously detect contradictions; if this fails, we explicitly provide the contradiction plan to assess genuine unsolvability. Only instances passing verification are retained.
@@ 3.2 Logic Puzzle Generation
Hitori. We formulate Hitori as a Constraint Satisfaction Problem and, using python-constraint, generate puzzles with unique solutions. For unsolvable instances, we inject conflicting constraints, for example forcing a row or column to contain duplicate numbers that cannot be removed without violating adjacency rules.
@@
Test Total 358 341 699
Table 1: Unified vertical statistics for UnsolvableQA train/test partitions. Train Maze has no difficulty split; AIME year ranges differ. (Note: AIME year ranges: train 1983–2023; test 2024–2025.)
@@ 3.3 Dataset Statistics and Quality Control
We construct UnsolvableQA to have approximately balanced solvable and unsolvable instances per domain, with minor residual imbalances caused by generator yields and verification constraints. Table 1 therefore reports aggregated statistics across the train and test partitions of UnsolvableQA. The test set contains 699 problems (358 solvable, 341 unsolvable), and the train set contains 637 problems (348 solvable, 289 unsolvable). The train partition pools all Maze instances without difficulty splits, whereas the test partition preserves Maze Easy. The AIME partitions cover different year ranges. Test prompts are diversified and paraphrased, while train prompts follow templated formats for scalable generation. All instances pass automated quality-control checks: logic puzzles are verified with multiple SAT/CSP/graph solvers, maze and programmatic tasks with deterministic solvers, and mathematically constructed problems with multi-pass LLM-based verification.
@@ 4.1 Problem Formulation
Figure 3: UnsolvableRL Reward Framework. We decompose the alignment into two orthogonal components: (a) Objective Detection, which incentivizes the model to distinguish between solvable and unsolvable queries. Crucially, we impose a penalty (-0.5) for falsely labeling solvable problems as unsolvable to safeguard the model’s reasoning drive. (b) Subjective Calibration, which dynamically regulates refusal behavior. By comparing the target threshold ($\tau$) with the model’s rollout accuracy ($\beta$), the mechanism encourages refusal on hard tasks (where $\tau>\beta$) and suppresses it on easy ones (where $\tau<\beta$) via the dynamic reward term $\lambda(\tau_{i}-\beta_{i})$.
@@ 4.3 Reward Design
This component evaluates the correctness of the model’s judgments on verifiably solvable instances. (1) A Correct Response (+1+1) is assigned when the model outputs the correct answer for a solvable instance. (2) For Incorrect Reasoning (0) on solvable instances, such as producing an incorrect final answer despite performing the calculation, the model receives a neutral reward. This asymmetric scheme emphasizes reasoning capability by rewarding successful problem solving while not explicitly penalizing unsuccessful but genuine reasoning attempts on solvable tasks.

### v2

@@ Abstract
Ensuring large language model (LLM) reliability requires distinguishing objective unsolvability (inherent contradictions) from subjective capability limitations (tasks exceeding model competence). Current LLMs often conflate these dimensions, leading to hallucinations in which they return confident answers to inherently unsolvable queries. To address this issue, we propose a multi-domain dataset containing both solvable and unsolvable questions, UnsolvableQA, together with an alignment framework, UnsolvableRL. First, we construct UnsolvableQA by “Reverse Construction” that systematically injects logical contradictions into otherwise valid reasoning chains. Second, we introduce UnsolvableRL, a reinforcement learning paradigm that balances objective unsolvability detection with calibrated confidence under capability limits. Empirically, our approach achieves near-perfect unsolvability detection ($90\%$ detection rate) and boosts solvable reasoning accuracy from 43.4% to 69.4% on Qwen3-4B-Instruct. Crucially, we identify a data–training interaction: strict alignment constraints induce Capability Collapse without unsolvable data, but act as a regularizer for rigor when such data are included, thereby improving overall robustness. Our code and data are available at https://github.com/sfasfaffa/unsolvableQA.
@@ 1 Introduction
As Large Language Models (LLMs) achieve strong performance on complex reasoning tasks (chen2025towards; wei2022chain; guo2025deepseek; jaech2024openai), a key research direction (qin2025largelanguagemodelsmeet) is ensuring they recognize their own limitations and behave reliably. Earlier work on confidence calibration (xiong2023can; li2024confidence; yoon2025reasoning; fisch2022calibrated) during reasoning underscores the need for models to know when not to answer. Further, studies on LLMs’ capability boundaries (chen2024unlocking) examine the self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; chen2025aware). More recently, TruthRL (wei2025truthrl) has encouraged LLMs to refuse when uncertain through a ternary reward mechanism.
@@
Figure 1: Comparison of alignment paradigms. We treat the solution space (solvable vs. unsolvable) and question difficulty (easy vs. hard) as orthogonal dimensions, requiring models to objectively detect solvability and subjectively calibrate confidence.
@@
However, a key limitation of these approaches is that they mainly focus on refusing questions that are solvable but currently beyond the model’s capabilities (the “Need to Calibrate” region in Figure 1 (a)). This emphasis overlooks a complementary, and more basic, challenge: identifying problems that are inherently unsolvable because they contain logical contradictions. More problematically, existing LLMs often misinterpret such unsolvability as high difficulty and may hallucinate answers even when the contradiction is explicit.
@@
To address this, our work treats Solution Space (solvable vs. unsolvable) and Question Difficulty (easy vs. hard) as orthogonal dimensions (Figure 1 (b)). Concretely, we argue that a reliable AI system must exhibit two behaviors: (1) Objective Detection, identifying internally contradictory problems and flagging them as “unsolvable” (yellow area); and (2) Subjective Calibration, safely refusing theoretically solvable but “difficult” problems that exceed the model’s reasoning capability. Without this distinction, models may hallucinate, forcing answers for inherently unsolvable problems.
@@
First, to support rigorous analysis of unsolvability, we develop a fully automated synthesis pipeline and introduce UnsolvableQA, a dataset of paired solvable and unsolvable instances spanning puzzles and mathematics. For programmatically verifiable puzzle tasks, we construct provably unsolvable instances by injecting code-level logical contradictions. For mathematics problems without reliable programmatic verification, inspired by lightman2024lets; ling2023deductive, we propose a “Reverse Construction” method: starting from a solvable problem and its correct rationale, we introduce a controlled contradiction into the rationale and then prompt LLMs to generate an unsolvable problem that matches the resulting flawed reasoning.
Then, to train models that remain robust on both solvable and unsolvable inputs, we propose UnsolvableRL, a reinforcement-learning (RL) method based on GRPO (shao2024deepseekmath). Unlike prior approaches that conflate refusal types, we define 3 distinct RL objectives: (1) Inherent unsolvability rewards generating an explicit refusal tag when solvability is detected; (2) False unsolvability Penalty discourages erroneous refusals by penalizing labeling solvable problems as unsolvable, thereby requiring explicit solvability checks before refusal; (3) Capability calibration uses a mechanism that rewards refusal only when the problem difficulty exceeds the LLMs’ current capability.
Empirical results show a synergistic effect: UnsolvableRL elevates the unsolvable problem detection rate of Qwen3-4B from 40.3% to 90.7% (near-perfect) and boosts its solvable problem accuracy from 43.4% to 69.4%, achieving an overall score of 80.1%. Our ablation study reveals three key insights: (1) models cannot learn to detect unsolvable problems without explicit unsolvable training data; (2) penalizing false refusals without such data causes “capability collapse”, where models stop refusing altogether; (3) when paired with unsolvable data, the penalty acts as a regularizer, trading a small drop in detection for stronger reasoning accuracy, a Pareto improvement over baselines.
@@
  • We first propose that reasoning alignment should treat objective unsolvability and subjective capability as orthogonal axes for reliable reasoning with clear decision boundaries.
  • We introduce UnsolvableQA, a dataset built with a “Reverse Construction” pipeline that injects logical contradictions into valid rationales to create verifiable unsolvable instances.
  • We propose UnsolvableRL, an RL paradigm with dynamic confidence thresholds and penalties for incorrect refusals, enabling models to solve feasible tasks, detect solvability, and calibrate confidence on hard inputs.
  • Controlled ablation studies validate the necessity and synergy of our two key designs: training on unsolvable data and applying a calibrated refusal penalty. This combination drives a Pareto improvement, maintaining high accuracy on solvable tasks while robustly detecting unsolvable ones.
@@ 2 Related Work
There are two primary concerns related to our work: automated dataset construction for reasoning and reinforcement learning approaches for incentivizing selective refusal.
Automatic Reasoning Dataset Construction.
As the reliance of Large Language Models on data quality increases, Automated Dataset Construction (ADC) has emerged as a prominent research direction (liu2024automatic; wang2023data). In the realm of reasoning, recent efforts on automatic data construction primarily focus on ensuring data solvability and verifiability through formal or programmatic means. Representative works such as Enigmata (chen2025enigmata) and BIGGSM (chen2024unlocking) have significantly enriched the resources for solvable reasoning tasks by employing synthetic puzzles and difficulty-graded annotations, respectively. Concurrently, research on verifying the correctness of rationale traces has also been advancing (ling2023deductive; chen2025towards).
In contrast, progress in dataset construction regarding problem unsolvability has been relatively slow. Earlier, SQuAD 2.0 (rajpurkar2018knowdontknowunanswerable) examined whether questions are answerable given a long context, but it does not adequately evaluate refusal behavior in complex logical-reasoning settings. Consequently, for reasoning tasks, existing data-synthesis efforts focus mainly on solvable instances while largely neglecting unsolvable questions, leaving a significant research gap.
@@ 3 UnsolvableQA Construction
Existing benchmarks mainly target solvable tasks, so models remain prone to hallucinations when assessing feasibility. To address this, we introduce UnsolvableQA, a dataset for detecting intrinsic contradictions in problem statements, using a dual-track methodology (in Figure 2) aligned with domain-specific verification constraints.
@@ 3.1 Mathematics Problem Generation via Reverse Construction
Data Generation Process. First, we validate each seed problem by confirming that a strong reasoning model (DeepSeek-Reasoner) can correctly solve it and produce a complete chain-of-thought, ensuring sufficient reasoning capability. Then, we employ DeepSeek-Reasoner to design logical contradictions via two complementary strategies: Constraint Contradiction, which introduces conflicting external conditions, and Axiom Contradiction, which modifies reasoning steps to violate fundamental mathematical axioms. The model generates a structured plan specifying the location and mechanism of contradiction injection.
Validation Process. Based on this plan, LLMs then reverse-engineers a new problem statement that preserves the stylistic and complexity characteristics of the seed problem while embedding the designed contradiction. Finally, we implement a two-tier verification protocol: the generated problem is first presented to DeepSeek-R1 to autonomously detect contradictions; if model fails to independently identify the contradiction, we proceed to the second tier. In this stage, we explicitly provide the model with both the final unsolvable problem statement and the predefined contradiction plan, thereby supplementing it with the intended contradiction information beyond the problem text alone. The model is then prompted to reassess the problem’s solvability given this complete contextual guidance. Only instances that are ultimately verified as unsolvable through this two-tier procedure are retained.
@@ 3.2 Logic Puzzle Generation
Hitori. We formulate Hitori as a constraint satisfaction problem and, using python-constraint to generate puzzles with unique solutions. For unsolvable instances, we inject conflicting constraints, for example forcing a row or column to contain duplicate numbers that cannot be removed without violating adjacency rules.
@@
Test Total 458 441 899
Table 1: Unified vertical statistics for UnsolvableQA train/test partitions. Train Maze has no difficulty split; AIME year ranges differ.
@@ 3.3 Dataset Statistics and Quality Control
We construct UnsolvableQA to maintain approximately balanced instances per domain (Table 1). The dataset comprises 899 test problems (458 solvable, 441 unsolvable) and 637 training problems (348 solvable, 289 unsolvable). While training prompts follow templated formats and pool Maze difficulties, test prompts are diversified and preserve fine-grained splits (including distinct AIME years [^1] AIME year ranges: train 1983–2023; test 2024–2025.). All instances undergo rigorous quality control: logic puzzles are verified via SAT/CSP/graph solvers, mazes via deterministic search, and math problems through multi-pass LLM verification. We also introduce two OOD testsets in Appendix A.2.
@@ 4 UnsolvableRL Methodology
Figure 3: UnsolvableRL framework with two orthogonal components: (a) Unsolvability Detection, which penalizes labeling solvable queries as unsolvable to preserve reasoning incentives. (b) Capability Calibration, which regulates refusal by comparing the target threshold ($\tau$) with rollout accuracy ($\beta$); it encourages refusal when $\tau>\beta$ and discourages it when $\tau<\beta$ via $\lambda(\tau_{i}-\beta_{i})$.
@@ 4.3 Reward Design
SOTA Baselines Qwen3-1.7B Qwen3-4B
Dataset Metric Deepseek GPT-5.1 Gemini-3 Instruct + UnsolRL Instruct + UnsolRL
Game24 S / U 98.0 / 77.5 14.0 / 50.0 94.0 / 94.0 78.0 / 23.0 84.0 / 100.0 81.0 / 17.0 94.5 / 99.0
Mean 87.8 32.0 94.0 50.5 92.0 49.0 96.8
HamCycle S / U 83.5 / 94.0 8.3 / 82.0 100.0 / 98.0 22.9 / 28.0 20.8 / 88.0 37.5 / 57.0 41.1 / 94.5
Mean 88.8 45.1 99.0 25.5 54.4 47.3 67.8
HamPath S / U 84.0 / 96.0 50.0 / 86.0 100.0 / 90.0 14.0 / 33.0 27.0 / 85.0 37.0 / 56.5 55.0 / 96.5
Mean 90.0 68.0 95.0 23.5 56.0 46.8 75.8
Hitori S / U 70.0 / 86.0 6.0 / 12.0 98.0 / 100.0 8.0 / 21.0 11.0 / 59.0 34.5 / 6.5 63.5 / 94.5
Mean 78.0 9.0 99.0 14.5 35.0 20.5 79.0
Maze (Easy) S / U 100.0 / 98.9 95.0 / 100.0 99.0 / 94.6 0.0 / 85.6 0.0 / 95.2 32.7 / 55.6 96.5 / 98.9
Mean 99.5 97.5 96.8 42.8 47.6 44.1 97.7
Maze (Hard) S / U 98.0 / 98.0 86.0 / 94.0 96.0 / 91.0 0.0 / 80.0 0.0 / 79.0 13.3 / 64.5 65.8 / 89.0
Mean 98.0 90.0 93.5 40.0 39.5 38.9 77.9
AIME 24-25 S / U 85.0 / 63.8 61.7 / 48.9 95.0 / 34.1 38.3 / 33.5 35.4 / 52.7 67.9 / 25.3 69.6 / 62.2
Mean 74.4 55.3 64.5 35.9 44.1 46.6 65.9
Overall S / U 88.4 / 87.7 45.9 / 67.6 97.4 / 86.0 23.0 / 43.4 25.5 / 79.8 43.4 / 40.3 69.4 / 90.7
Mean 88.1 56.7 91.7 33.2 52.7 41.9 80.1
Table 2: Detailed performance comparison. For each dataset, the first row reports Solvable Accuracy (S) / Unsolvable Detection Rate (U), and the second row reports their Mean (M). We bold the global best score across all models for each metric. Qwen3-4B + UnsolRL achieves SOTA performance in Unsolvable Detection (Overall U) and Game24 Mean accuracy.
@@
It evaluates the correctness of LLMs’ judgments on verifiably solvable instances. (1) A Correct Response (+1+1) is assigned when the model outputs the correct answer for a solvable instance. (2) For Incorrect Reasoning (0) on solvable instances, such as producing an incorrect final answer despite performing the calculation, the model receives a neutral reward. This asymmetric scheme emphasizes reasoning capability by rewarding successful problem solving while not explicitly penalizing unsuccessful but genuine reasoning attempts on solvable tasks.